A 950lb horse requires treatment with metronidazole at a dose of 15mg/Kg t.i.d. Metronidazole is available as 500 mg tablets. Calculate the total number of tablets for 5 days
Answer #1

Weight of horse is 950lb. This is approximately 432kgs. (1Kg=2.2lbs)

Required dose of metronidazole is 15mg/Kg t.i.d. This means 15mg/Kg three times a day.

Total dosage of metronidazole to be given in one day is = weigh in kilograms * dosage per kilogram * 3

which is 432*15*3 = 19,440 grams.

Total number of days is 5.

dosage required for 5 days = dosage required in one day * number of days

which is 19,440*5= 97,200 grams.

Available tablets of metronidazole contains 500mg.

Number of tablets required for 5 days = total dosage required / amount of drug per single tablet

which is 97,200/500 = 194.4 which is 195 tablets approximately

number of tablets per day is = tablets required for 5 days / 5

which is 195/5 = 39.

The total number of tablets for 5 days is 195.

This is a brief discussion on the given topic. Hope this helps you out. Please feel free to give a like. Contact us if you need further detail explanation. Thank you and have a great day.